1/2x+ 2y=6
X-6y=2

Answer by TimothyLamb(4379)

---
1/2x + 2y = 6
x - 6y = 2
---
put the system of linear equations into standard form
---
0.5x + 2y = 6
x - 6y = 2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 8
y= 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php